OpenStudy (anonymous):

Solve A = 2p – 3r for r.

OpenStudy (amistre64):

r aint that hard to get at; what would you try?

OpenStudy (angela210793):

1st: -3r=A-2p 3r=2p-A r=(2p-A)/3

OpenStudy (anonymous):

first of all A-2p=-3r solve for r from there

OpenStudy (amistre64):

id try dividing at all by -3 first :P

OpenStudy (anonymous):

A-2p/-3=r

OpenStudy (anonymous):

r = r = r = A – 2p + 3 r = A – 2p

OpenStudy (anonymous):

no one of thoose is the answer